# Design and Performance of a Universal SiPM Readout System for X- and Gamma-Ray Missions

### v1

@@ 4.2 Non-linear Gain
When calculating the dynamic range in the previous section, we assumed a fully linear relation between keV and ADC. This is, however, not correct due to saturation of the amplifiers in the Citiroc 1A ASIC. This has previously been discussed in Kole et al. (2024). This saturation can be noticed in a spectrum which is not fully contained within the dynamic range, or when plotting the high gain ADC value vs. the low gain ADC value. This is shown in figure 8 which shows the (non-pedestal subtracted) high gain (top left) and low gain (top right) for a second channel from the same measurement as used for figure 7. Whereas in figure 7 the photo-peak is fully contained, here the photo-peak can be clearly seen to be cut off in the high gain spectrum which shows no events beyond 3800 ADC whereas the ADC range extends up to 4095 ADC. In the low gain the peak is fully visible, while, in addition, cosmic ray induced events can be seen up to 4095 ADC. When plotting the high gain against the low gain ADC values for this measurement, the curve shown in the bottom panel of figure 8 is achieved. The high gain ADC value is linear with the low gain, and therefore with the keV input, up to approximately 2000 ADC. Above this, the amplifier in the ASIC starts to saturate, until at 3100 ADC it is fully saturated. As a result, no ADC values exceeding this value are observed. This indicates that the saturation of the amplifier causes $25\%$ of the dynamic range in the high gain to be lost. However, above this range one can rely on the low gain ADC values where no such saturation is observed.
It should further be noted that for the channel studied here, the dynamic range significantly differs from that of the channel studied in the previous section. This is a result of a difference in the gain of the two channels. Both channels are readout with SiPM channels which are operated at equal over-voltages [^2] The breakdown voltages of the various SiPM channels and their temperature dependence were calibrated prior to mounting them on the PCB. We do, however, observe that for the channel used here we have a gain of approximately 49 ADC/p.e., whereas for the channel of figure 7 the gain is 25 ADC/p.e.. This gain can be adjusted using the 6 bit DAC of the high gain which was here set to be equal for all channels. In case a higher uniformity between the various channels is desired, this DAC value can be used to equalize the gains.
@@ 4.3 Photo-electron Resolution
Both spectra of figures 7 and 8 show numerous photo-electron peaks, also called fingers in the field of SiPMs. The capability to observe and fit numerous fingers allows for the calibration of the gain of the channels, thereby allowing to accurately produce an instrument response. In addition, by fitting and adjusting the gain, one can produce a uniform sensitivity in all the detector channels. We can test whether these fingers show their expected behaviour. Each finger in the spectrum corresponds to one fired microcell in the SiPM, therefore the distance between these fingers should be constant. In addition, the width of the fingers increases with their number (De Angelis, 2019) according to:
@@
where $\sigma_{p.e}$ is the width of the finger retrieved from a Gaussian fit, $N_{p.e.}$ is the number of the finger and $\sigma_{e}$ is the width of the pedestal electronic noise. The latter contains the width of the electronics which can be measured using the width of the pedestal as presented in section 3.2. However, whereas the pedestal only contains the noise induced by the electronics, the zeroth finger also contains electronic noise from the SiPM itself, causing $\sigma_{e}$ to be larger than the noise measured through the pedestal. $\sigma_{1}$ represents the increase in the width due to amplification of the noise. Using this relation for the width of the fingers, free parameters for the height of the various fingers, and using a single fitting parameter for the width between peaks, we can fit the first 12 fingers visible in the high gain spectrum of Figure 8. The result can be seen in figure 9, i.e. the fingers follow the expected behaviour with a fixed distance of 49 ADC, a $\sigma_{e}=11.9$ ADC and $\sigma_{1}=3.4$ ADC.
@@
5 Performance with GAGG and MPPC S13661
@@ 5.1 Test Setup
The SiPM readout system was additionally tested for use with a high-Z scintillator. To date, the system has been tested using both CeBr3 and GAGG:Ce. For CeBr3 only basic tests were performed which showed that the scintillator can be read out successfully. The combination of the system together with an array of 64 GAGG crystals was tested in more detail using the LARIX-A X-ray facility of the University of Ferrara (Virgilli et al., 2013). Similar to the tests with plastic scintillators described in section 4, the system was equipped with 4 arrays of $4\times 4$ MPPCs from Hamamatsu of the S13660 type, the S13661-6075PE-04. The array of GAGG scintillators consists of crystals with dimensions of $6\times 6\times 15mm^{3}$ which are placed in a Barium sulfate (BaSO4) housing which provides optical decoupling of the crystals in the array as well as mechanical support. The crystals are arranged such that they couple exactly to the SiPM array placed on the readout system. Finally, the GAGG array is mechanically coupled to the readout system using a custom-made carbon fiber housing, which ensures a light-tight readout system as well as mechanical stability.
@@ 5.3 Energy Resolution
Using the various features in these measured spectra we can look at the energy resolution, defined here as the full-width half maximum (FWHM) over the mean, as a function of energy. This can be measured both using the high gain, for energies up to 40 keV above which the gain becomes non-linear, as well as using the low gain starting from 20 keV at which point the absorption peaks are significantly separated from the noise region. The results, based on fits of both photo-peaks and escape peaks, can be seen in figure 11 which shows the resolution measured in low gain on the right and high gain on the left. It can be seen that at low energies the energy resolution in the high gain is significantly better than that measured in the low gain, indicating deterioration of the energy resolution at low ADC values in the low gain. Around 40 keV the two resolution values become comparable. The K-edge from Gallium, at 50 keV, starts to become clearly visible. Above the K-edge it becomes significantly more likely that the energy is deposited by two individual interactions, one from the K-edge absorption and one from the remaining photon energy, rather than one. This deteriorates the energy resolution. The energy resolution achieved here, which was measured without optimizing the setup for this purpose, is comparable to those mentioned in for example Yoneyama et al. (2018), where an energy resolution of $\approx 45\%$ is presented for 22 keV and $\approx 25\%$ is presented for 88 keV. As will be discussed next in section 5.4, the optical coupling between the GAGG and the SiPM can likely be improved significantly, as here no optical coupling was used. This would result in more optical photons/keV reaching the SiPM and therefore in a better energy resolution.
@@ 5.5 Further potential optimization
The 640 keV calculated here assumes a fully linear relation between the input energy and the measured ADC value. We know already, as discussed in section 4.2, that for the high gain the amplifier is not fully linear at high ADC values resulting in some saturation. In addition, saturation of the number of microcells in a SiPM array will also cause non-linearity, thereby limiting the dynamic range. We can calculate the effect of saturation of the used SiPM type, the S13361-6075, as well as that for the versions with $50\mu m$ and $25\mu m$ cells. For this we use the following relation (De Angelis et al., 2023):
@@ 6.2 Irradiation
One of the main concerns when using COTS components in the design is their resilience to radiation. From all the components used, only the Citiroc 1A ASICs are at TRL 8 (Weeroc, 2025), thereby indicating that they are radiation tolerant. Neither the IGLOO FPGA, the DC/DC converter used to power the SiPMs, or the ADCs are known to be radiation tolerant. Of particular concern was the DC/DC converter, as initially the C11204-01 by Hamamatsu was considered for use. This DC/DC converter, which is specifically designed for powering SiPMs, failed a basic proton irradiation test resulting in the change to the T3482 DC/DC in the current design.
@@
It should of course be noted that the SiPMs are known to be prone to radiation damage. For the POLAR-2 mission these were therefore subject to a detailed dedicated set of tests which are presented in De Angelis et al. (2023).
@@ 6.3 Vibration and Shock Testing
The parameters used for the sinusoidal vibrations, applied to all 3 axes, are summarized in table 2. Similarly, those for the random vibrations are summarized in table 3. Finally, those used for the shocks are shown in table 4. For the shock tests, the levels were incremented in small steps until the desired levels were measured using sensors placed on the setup. The vibration and shock levels applied here were derived from those applied to the POLAR mission during its final space qualification tests. As the POLAR detector was launched while mounted on the Tiangong-2 spacelab, these levels are relatively high compared to a normal space launch (Produit et al., 2018).
@@
Figure 16: Definition of the X-Y-Z frame with respect to system. The Z axis is defined as the vertical direction (along the scintillator length), and the X and Y directions are defined in the horizontal plane along the scintillators. The readout system is placed in the aluminium frame. Accelerometers are placed on the flange of the module (green) and near the dampers (red) in order to monitor the resonance spectrum during and between the different tests (taken from De Angelis (2023), with permission).
@@ 7 Conclusion
A readout system for 64 SiPM channels was developed. The system contains only COTS components, allowing to have a cost of several kUSD (approximately 3000 USD when developing 10 units), while also avoiding many of the common export problems found with such electronic systems for space missions. The system furthermore provides a large autonomous readout for 64 SiPM channels with a total power consumption of approximately 1.8 W.
@@
As such, the system can be used both for detectors like POLAR-2 which use only plastic scintillators, as well as spectrometers based on high-Z scintillators. Also, the system can be used to read out combinations of these two scintillators, which can be of use for future polarimeters like SPHiNX (Pearce et al., 2019), or the upcoming evolution of LEAP proposed by some of the authors as an upcoming NASA mission.
@@
Furthermore, the system will be tested using linearly graded SiPMs (Gola et al., 2020). Such SiPMs have several (4 or 6) anodes for each channel, where the relative output of the various anodes provides information on the position of the fired microcell. As such, LG-SiPMs can be used to produce large-scale position sensitive detectors, or single LG-SiPM channels can be used to read out many small scintillators such as fibers. The system described here would be capable of reading out up to 16 LG-SiPMs (of the 4 anode readout type), thereby allowing to read out potentially hundreds of scintillating fibers using only 1.8 W.

### v2

@@ 4.2 Non-linear Gain
When calculating the dynamic range in the previous section, we assumed a fully linear relation between keV and ADC. This is, however, not correct due to saturation of the amplifiers in the Citiroc 1A ASIC. This has previously been discussed in [31]. This saturation can be noticed in a spectrum which is not fully contained within the dynamic range, or when using the direct charge injection to produce a figure such as figure 4. Figure 9 shows the (non-pedestal subtracted) high gain (top left) and low gain (top right) for a second channel from the same measurement as used for figure 8. Whereas in figure 8 the photo-peak is fully contained, here the photo-peak can be clearly seen to be cut off in the high gain spectrum which shows no events beyond 3800 ADC whereas the ADC range extends up to 4095 ADC. In the low gain the peak is fully visible, while, in addition, cosmic ray induced events can be seen up to 4095 ADC. When plotting the high gain against the low gain ADC values for this measurement, the curve shown in the bottom panel of figure 9 is achieved. The high gain ADC value is linear with the low gain, and therefore with the keV input, up to approximately 2000 ADC. Above this, the amplifier in the ASIC starts to saturate, until at 3100 ADC it is fully saturated. As a result, no ADC values exceeding this value are observed. This indicates that the saturation of the amplifier causes $25\%$ of the dynamic range in the high gain to be lost. However, above this range one can rely on the low gain ADC values where no such saturation is observed.
It should further be noted that for the channel studied here, the dynamic range significantly differs from that of the channel studied in the previous section. This is a result of a difference in the gain of the two channels. Both channels are readout with SiPM channels which are operated at equal over-voltages [^5] The breakdown voltages of the various SiPM channels and their temperature dependence were calibrated prior to mounting them on the PCB. We do, however, observe that for the channel used here we have a gain of approximately 49 ADC/p.e., whereas for the channel of figure 8 the gain is 25 ADC/p.e.. This gain can be adjusted using the 6 bit DAC of the high gain which was here set to be equal for all channels. In case a higher uniformity between the various channels is desired, this DAC value can be used to equalize the gains. This has previously been achieved for POLAR-2 and a detailed discussion of this is presented in [25]. For further information, which requires details on the calibration of POLAR-2: the full calibration of one POLAR-2 module, and the resulting distribution of the gain from various channels is discussed in detail in [25].
@@ 4.3 Photo-electron Resolution
Both the spectra of figures 8 and 9 show numerous photo-electron peaks, also called fingers in the field of SiPMs. The capability to observe and fit numerous fingers allows to measure the number of photo-electrons produced per deposited keV and therefore study the photo-detection efficiency of each channel. Subsequently, one can fine tune the bias voltages for each channel, as well as the electronic gain to produce a uniform sensitivity in all the detector channels. Within one channel, each finger in the spectrum corresponds to one photo-electron and therefore on fired microcell in the SiPM. The distance between the fingers should be constant. In addition, the width of the fingers increases with their number [33] according to:
@@
where $\sigma_{p.e}$ is the width of the finger retrieved from a Gaussian fit, $N_{p.e.}$ is the number of the finger and $\sigma_{e}$ is the width of the pedestal electronic noise. The latter contains the width of the electronics induced noise which can be measured using the width of the pedestal as presented in section 3.3. However, whereas the pedestal only contains the noise induced by the electronics, the zeroth finger also contains electronic noise from the SiPM itself, causing $\sigma_{e}$ to be larger than the noise measured through the pedestal. $\sigma_{1}$ represents the increase in the width due to amplification of the noise. Using this relation for the width of the fingers, leaving the height of the various fingers as a free parameter, and using a single fitting parameter for the width between peaks, we can fit the first 12 fingers visible in the high gain spectrum of Figure 9. The result can be seen in figure 10, i.e. the fingers follow the expected behaviour with a fixed distance of 49 ADC, a $\sigma_{e}=11.9$ ADC and $\sigma_{1}=3.4$ ADC. Using the position of the photo-peak, we know 49 ADC corresponds to approximately 1 keV. Therefore we know that $\sigma_{e}=0.24$ keV and $\sigma_{1}=0.07$ keV. The ratio of $\sigma$ over the finger separation observed here is primarily a result of the choice of SiPM. Using SiPMs with a smaller terminal capacitance such as those with smaller channels, rather than the $6\times 6\,\mathrm{mm^{2}}$ ones used here, will result in smaller widths. Initial tests of the same front-end with Hamamatsu SiPMs of the S14 type deemed to be unsuccessful as this type has a larger terminal capacitance, thereby making it almost impossible to see the fingers in the spectrum [25].
@@
5 Performance with GAGG and MPPC S13361
@@ 5.1 Test Setup
The SiPM readout system was additionally tested for use with a high-Z scintillator. To date, the system has been tested using both CeBr3 and GAGG:Ce. For CeBr3 only basic tests were performed which showed that the scintillator can be read out successfully. The combination of the system together with an array of 64 GAGG crystals was tested in more detail using the LARIX-A X-ray facility of the University of Ferrara [34]. Similar to the tests with plastic scintillators described in section 4, the system was equipped with 4 arrays of $4\times 4$ MPPCs from Hamamatsu of the S13360 type, the S13361-6075PE-04. The array of GAGG scintillators consists of crystals with dimensions of $6\times 6\times 15\,\mathrm{mm}^{3}$ which are placed in a barium sulfate (BaSO4) housing which provides optical decoupling of the crystals in the array as well as mechanical support. Figure 11 shows the array, which is coupled directly to the SiPM array shown in figure 2. A 1 mm thick carbon fiber cover was placed over the array to provide light tightness and to couple it mechanically to the frame. The combination was placed such that the collimated beam is perpendicular to the array.
Figure 11: The GAGG array used to study the behavior of the system at the LARIX-A X-ray facility. The GAGG crystals match the $6\times 6\,\mathrm{mm^{2}}$ area of the SiPMs and have a height of $15\,\mathrm{mm}$. They are separated by a thin layer (white) of barium sulfate.
The crystals are arranged such that they couple exactly to the SiPM array placed on the readout system. Finally, the GAGG array is mechanically coupled to the readout system using a custom-made carbon fiber housing, which ensures a light-tight readout system as well as mechanical stability.
@@ 5.3 Energy Resolution
Using the various features in these measured spectra we can look at the energy resolution, defined here as the full-width half maximum (FWHM) over the mean, as a function of energy. This can be measured both using the high gain, for energies up to 40 keV above which the gain becomes non-linear, as well as using the low gain starting from 20 keV at which point the absorption peaks are significantly separated from the noise region. The results, based on fits of both photo-peaks and escape peaks, can be seen in figure 13 which shows the resolution measured in low gain on the right and high gain on the left. It can be seen that at low energies the energy resolution in the high gain is significantly better than that measured in the low gain, indicating deterioration of the energy resolution at low ADC values in the low gain. Around 40 keV the two resolution values become comparable. The K-edge from Gallium, at 50 keV, starts to become clearly visible. Above the K-edge it becomes significantly more likely that the energy is deposited by two individual interactions, one from the K-edge absorption and one from the remaining photon energy, rather than one. This deteriorates the energy resolution. The energy resolution achieved here, which was measured without optimizing the setup for this purpose, is comparable to those mentioned in for example [35], where an energy resolution of $\approx 45\%$ is presented for 22 keV and $\approx 25\%$ is presented for 88 keV. As will be discussed next in section 5.4, the optical coupling between the GAGG and the SiPM can likely be improved significantly, as here no optical coupling was used. This would result in more optical photons/keV reaching the SiPM and therefore in a better energy resolution.
Although the energy resolution looks good at the lower energies tested here, it would be useful to also irradiate this setup with higher energy photons. For this detailed tests with both radioactive sources as well as synchrotron beams with energies up to $500\,\mathrm{keV}$ will be performed in the future.
@@ 5.5 Further potential optimization
The 640 keV calculated here assumes a fully linear relation between the input energy and the measured ADC value. We know already, as discussed in section 4.2, that for the high gain the amplifier is not fully linear at high ADC values resulting in some saturation. In addition, saturation of the number of microcells in a SiPM array will also cause non-linearity, thereby limiting the dynamic range. We can calculate the effect of saturation of the used SiPM type, the S13361-6075, as well as that for the versions with $50\mu m$ and $25\mu m$ cells. For this we use the following relation [24]:
@@ 6.2 Irradiation
One of the main concerns when using COTS components in the design is their resilience to radiation. From all the components used, only the Citiroc 1A ASICs are at TRL 8 [36], thereby indicating that they are radiation tolerant. Neither the IGLOO FPGA, the DC/DC converter used to power the SiPMs, or the ADCs are known to be radiation tolerant. Of particular concern was the DC/DC converter, as initially the C11204-01 by Hamamatsu was considered for use. This DC/DC converter, which is specifically designed for powering SiPMs, failed a basic proton irradiation test resulting in the change to the T3482 DC/DC in the current design.
@@
It should of course be noted that the SiPMs are known to be prone to radiation damage. For the POLAR-2 mission these were therefore subject to a detailed dedicated set of tests which are presented in [24].
@@ 6.3 Vibration and Shock Testing
The parameters used for the sinusoidal vibrations, applied to all 3 axes, are summarized in table 3. Similarly, those for the random vibrations are summarized in table 4. Finally, those used for the shocks are shown in table 5. For the shock tests, the levels were incremented in small steps until the desired levels were measured using sensors placed on the setup. The vibration and shock levels applied here were derived from those applied to the POLAR mission during its final space qualification tests. As the POLAR detector was launched while mounted on the Tiangong-2 spacelab, these levels are relatively high compared to a normal space launch [4].
@@
Figure 18: Definition of the X-Y-Z frame with respect to system. The Z axis is defined as the vertical direction (along the scintillator length), and the X and Y directions are defined in the horizontal plane along the scintillators. The readout system is placed in the aluminium frame. Accelerometers are placed on the flange of the module (in green) and near the dampers (in red) in order to monitor the resonance spectrum during and between the different tests (taken from [25], with permission).
@@ 7 Conclusion
A readout system for 64 SiPM channels was developed. The system contains mostly COTS components, allowing to have a cost of several kUSD (approximately 3000 USD when developing 10 units), while also avoiding many of the common export problems found with such electronic systems for space missions. The system furthermore provides a large autonomous readout for 64 SiPM channels with a total power consumption of approximately 1.8 W.
@@
As such, the system can be used both for detectors like POLAR-2 which use only plastic scintillators, as well as spectrometers based on high-Z scintillators. Also, the system can be used to read out combinations of these two scintillators, which can be of use for future polarimeters like SPHiNX [37], or the upcoming evolution of LEAP proposed by some of the authors as an upcoming NASA mission.
@@
Furthermore, the system will be tested using linearly graded SiPMs [23]. Such SiPMs have several (4 or 6) anodes for each channel, where the relative output of the various anodes provides information on the position of the fired microcell. As such, LG-SiPMs can be used to produce large-scale position sensitive detectors, or single LG-SiPM channels can be used to read out many small scintillators such as fibers. The system described here would be capable of reading out up to 16 LG-SiPMs (of the 4 anode readout type), thereby allowing to read out potentially hundreds of scintillating fibers using only 1.8 W.